Three times the greater of two consecutive odd integers is five less than four times the smaller. Find the two numbers.
Answer by rapaljer(4671)

Let x = first odd integer (smaller)
x+2 = second odd integer (larger)

3(x+2) = 4x-5
3x+ 6 = 4x -5

Subtract 3x from each side:
3x-3x + 6 = 4x-3x -5
6=x-5

Add +5 to each side:
6+5= x-5+5
11=x
x+2= 13.

Smaller = 11, Larger = 13.

Check:
Three times the larger is five less than 4 times the smaller.
3(13)= 4(11) -5
39= 44-5

It checks!!

For additional explanation on Word Problems, in particular "Consecutive Integer" problems, please see my own website. Do a "Bing" or "Google" search for my last name "Rapalje". Look for "Rapalje Homepage" near the top of these searches. Then near the top of my Homepage, look for "Basic, Intermediate and College Algebra". Choose "Basic Algebra", "Chapter 1", "Section 1.10 Word Problems Part I". See also the "MATH IN LIVING COLOR" pages that go with this section.

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus